# Mathematical artificial data for operator learning

### v1

@@ 1 Introduction
The remainder of this paper is organized as follows. In Section 2, we introduce the Mathematical Artificial Data (MAD) framework, detailing its theoretical foundation, data-generation procedures, and evaluation metrics. Section 3 presents numerical results, and Section 4 discusses the implications, limitations, and future directions.
@@ 2.2.2 Fundamental solution-based artificial data (MAD1)
Moreover, for the three-dimensional Laplace equation on $[0,1]^{3}$, we employed the 3D fundamental solution $-\frac{1}{4\pi\|x-y\|}$, and similarly sampled its linear combination coefficients from $\mathcal{N}(0,1)$.
@@ 2.2.4 Data sampling in PINN-based methods
[…]
@@ 2.3 Experimental setup
  • For source-free equations: An unstacked DeepONet was used with a branch network of size $[200,1000]$ and a trunk network configured as $[2,110,\allowbreak 110,110,110,1000]$. This configuration served as the baseline for the following modifications:
@@
For the 3D Laplace equation over $[0,1]^{3}$, we employ a DeepONet with:
@@
For the two MAD1-FNO variants on the 2D Laplace equation over $[0,1]^{2}$, the input tensor has shape $[C_{\rm in}\times N\times N]$, where $C_{\rm in}=2$ for MAD1-FNO1 or $4$ for MAD1-FNO2. The network consists of:
  • Input projection: 1×1 convolution mapping $C_{\rm in}\to 64$ channels, producing a tensor of shape $[64\times N\times N]$.
  • Four Fourier layers: each layer applies a spectral convolution on the lowest 32 modes per dimension with 64 input/output channels and zero-padding (output $[64\times N\times N]$), followed by a 1×1 convolution ($64\to 64$) and a GELU activation, defined as $\mathrm{GELU}(x)=x\,\Phi(x)$ with $\Phi$ the standard normal cumulative distribution function (CDF) [20].
  • Output head: two successive 1×1 convolutions mapping $64\to 64\to 1$, yielding the final output tensor of shape $[1\times N\times N]$.
@@ 3 Results
In this work, according to the mathematical properties of different types of equations, we propose three MAD variants—MAD0 for equations with source terms, MAD1 (fundamental-solution based) and MAD2 for source-free equations—and evaluate them through extensive numerical experiments on the Laplace, Poisson, and Helmholtz equations to assess their accuracy, efficiency, and generalization ability.
@@
Leveraging MAD’s rapid, exact analytical data generation—decoupled entirely from network design—our framework can produce large, high-quality datasets in seconds. Moreover, it integrates seamlessly with any modern neural architectures or optimization algorithms. In particular, for each method, we designed two test sets (test set 1 and test set 2) sampled using completely different schemes, thereby providing a more convincing comparison of generalization performance.
@@
Consequently, the finite-difference solver was run on the fine grid, $h=0.001$. Under this criterion, MAD produced the 200 Helmholtz samples in 2.19 s on an Intel Core i7-12800HX CPU, whereas the high-precision FD solver required 4,663.17 s, underscoring MAD’s overwhelming speed advantage.
@@ 3.1 General equations with source using MAD0
Table 1: Comparison of PINN-based methods and MAD0 for solving the Poisson (equation (4)) and Helmholtz (equation (5)) equations with source. Test Set 1 comprises analytically generated solutions closely aligned with the MAD0 training set, whereas Test Set 2 is created using the five-point central difference method on a $0.001\times 0.001$ grid over the $[0,1]\times[0,1]$ domain, ensuring high accuracy for evaluating the performance of different models.
@@ 3.2 Source-free equations using MAD1 (based on fundamental solutions)
Table 2: Comparison of PINN-based and MAD1 methods for solving the equations (4)-(5) with $f=0$. Each value in the table represents the mean over four independent experiments. Test Set 1 comprises analytically generated solutions closely aligned with the MAD1 training set, whereas Test Set 2 is created using the five-point central difference method on a $0.001\times 0.001$ grid over the $[0,1]\times[0,1]$ domain, ensuring high accuracy for evaluating the performance of different models.
@@ 3.5 Experiments using the FNO network architecture
As a demonstration outside of the DeepOnet framework, we also trained a standard Fourier neural operator (FNO) using MAD synthesis data of the 2D Laplace equation on $[0,1]^{2}$. Traditionally, such data are obtained by running high-fidelity numerical solvers on fine meshes, which incurs significant computational cost. In contrast, our MAD framework generates exact analytic solutions almost instantly.
@@ 3.6 A test of solving the 3D Laplace equation
We also evaluated the MAD1 framework combined with a DeepONet backbone on the 3D Laplace equation over $[0,1]^{3}$. After 10 000 training epochs, the model achieved a training loss of $1.61\times 10^{-5}$ and a relative $L_{2}$ error of $2.03\times 10^{-2}$. These results demonstrate that our MAD framework can be readily extended to higher-dimensional boundary-value problems.

### v2

@@ 1 Introduction
The remainder of this paper is organized as follows. In Section 2, we introduce the Mathematical Artificial Data (MAD) framework, detailing its theoretical foundation, data‐generation procedures, and evaluation metrics. Section 3 presents numerical results, and Section 4 discusses the implications, limitations, and future directions.
@@ 2.2.2 Fundamental solution-based artificial data (MAD1)
Moreover, for the three-dimensional Laplace equation on $0,1]^{3}$, we employed the 3D fundamental solution $\frac{1}{4\pi\|x-y\|}$, and similarly sampled its linear combination coefficients from $\mathcal{N}(0,1)$.
@@ 2.2.4 Data sampling in PINN-based methods
$f_{\text{smoothed}}(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\int_{\Omega}f_{\text{raw}}(y)\exp\left(-\frac{|x-y|^{2}}{2\sigma^{2}}\right)\,dy,$
@@ 2.3 Experimental setup
  • For source-free equations: An unstacked DeepONet was used with a branch network of size $[200,1000]$ and a trunk network configured as $2,110,\allowbreak 110,110,110,1000]$. This configuration served as the baseline for the following modifications:
@@
For the 3D Laplace equation over $0,1]^{3}$, we employ a DeepONet with:
@@
For the two MAD1-FNO variants on the 2D Laplace equation over $0,1]^{2}$, the input tensor has shape $C_{\rm in}\times N\times N]$, where $C_{\rm in}=2$ for MAD1-FNO1 or $4$ for MAD1-FNO2. The network consists of:
  • Input projection: 1×1 convolution mapping $C_{\rm in}\to 64$ channels, producing a tensor of shape $64\times N\times N]$.
  • Four Fourier layers: each layer applies a spectral convolution on the lowest 32 modes per dimension with 64 input/output channels and zero-padding (output $64\times N\times N]$), followed by a 1×1 convolution ($64\to 64$) and a GELU activation, defined as $\mathrm{GELU}(x)=x\,\Phi(x)$ with $\Phi$ the standard normal cumulative distribution function (CDF) [20].
  • Output head: two successive 1×1 convolutions mapping $64\to 64\to 1$, yielding the final output tensor of shape $1\times N\times N]$.
@@ 3 Results
In this work, according to the mathematical properties of different types of equations, we propose three MAD variants—MAD0 for equations with source terms, MAD1 (fundamental‐solution based) and MAD2 for source‐free equations—and evaluate them through extensive numerical experiments on the Laplace, Poisson, and Helmholtz equations to assess their accuracy, efficiency, and generalization ability.
@@
Leveraging MAD’s rapid, exact analytical data generation—decoupled entirely from network design—our framework can produce large, high‐quality datasets in seconds. Moreover, it integrates seamlessly with any modern neural architectures or optimization algorithms. In particular, for each method, we designed two test sets (test set 1 and test set 2) sampled using completely different schemes, thereby providing a more convincing comparison of generalization performance.
@@
Consequently, the finite‐difference solver was run on the fine grid, $h=0.001$. Under this criterion, MAD produced the 200 Helmholtz samples in 2.19 s on an Intel Core i7-12800HX CPU, whereas the high-precision FD solver required 4,663.17 s, underscoring MAD’s overwhelming speed advantage.
@@ 3.1 General equations with source using MAD0
Table 1: Comparison of PINN-based methods and MAD0 for solving the Poisson (equation (4)) and Helmholtz (equation (5)) equations with source. Test Set 1 comprises analytically generated solutions closely aligned with the MAD0 training set, whereas Test Set 2 is created using the five-point central difference method on a $0.001\times 0.001$ grid over the $0,1]\times[0,1]$ domain, ensuring high accuracy for evaluating the performance of different models.
@@ 3.2 Source-free equations using MAD1 (based on fundamental solutions)
Table 2: Comparison of PINN-based and MAD1 methods for solving the equations (4)-(5) with $f=0$. Each value in the table represents the mean over four independent experiments. Test Set 1 comprises analytically generated solutions closely aligned with the MAD1 training set, whereas Test Set 2 is created using the five-point central difference method on a $0.001\times 0.001$ grid over the $0,1]\times[0,1]$ domain, ensuring high accuracy for evaluating the performance of different models.
@@ 3.5 Experiments using the FNO network architecture
As a demonstration outside of the DeepOnet framework, we also trained a standard Fourier neural operator (FNO) using MAD synthesis data of the 2D Laplace equation on $0,1]^{2}$. Traditionally, such data are obtained by running high-fidelity numerical solvers on fine meshes, which incurs significant computational cost. In contrast, our MAD framework generates exact analytic solutions almost instantly.
@@ 3.6 A test of solving the 3D Laplace equation
We also evaluated the MAD1 framework combined with a DeepONet backbone on the 3D Laplace equation over $0,1]^{3}$. After 10 000 training epochs, the model achieved a training loss of $1.61\times 10^{-5}$ and a relative $L_{2}$ error of $2.03\times 10^{-2}$. These results demonstrate that our MAD framework can be readily extended to higher-dimensional boundary-value problems.